OpenStudy (anonymous):

help me solve please 2

OpenStudy (anonymous):

what have you tried so far?

OpenStudy (anonymous):

nothing im not sure how to work out these problems

OpenStudy (anonymous):

so for the first one do you have this?\[\large 2 < d + \frac{1}{2} < 4(\frac{1}{2})\]:

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but a negative in front of the 2

OpenStudy (anonymous):

so,\[\large -2 < d + \frac{1}{2} < 4(\frac{1}{2})\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So we have: \[\large -2 < d + \frac{1}{2} < 2\]

OpenStudy (anonymous):

now we minus 1/2 from everywhere to get: \[\large -2-\frac{1}{2} < d < 2-\frac{1}{2}\] and then we get (by minusing):\[\large -\frac{5}{2} < d < \frac{3}{2}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

oh yeah now i see

OpenStudy (anonymous):

is that al you do for that problem

OpenStudy (anonymous):

yip :) so do you understand?

OpenStudy (anonymous):

the other two are easier so if you understand this one you can most probably do the other two

OpenStudy (anonymous):

thank you mucho